# PP-TIL: Personalized Planning for Autonomous Driving with Instance-based Transfer Imitation Learning

### v1

@@ III-A Problem Formulation
We assume the current time step is $t$. The agent state at the current time is represented as $\bm{s}_{t}=[x_{t},y_{t},\theta_{t},v_{t}]$, where $(x_{t},y_{t})$ is two-dimensional space coordinate, $\theta_{t}$ is heading angle, and $v_{t}$ is velocity. The trajectory $\xi_{0:t}$ is represented as a set of discrete points $\{s_{0},...,s_{t}\}$. The model predicts the future trajectory of agent $i$ for the next $T$ time steps, and it’s represented as $\hat{\xi}^{i}_{t:t+T}$. The output of the neural network is the control action sequence and denoted as ${\Pi}^{m}_{t:t+T}(m={1,...,N_{m}})$, where $N_{m}$ is the number of multi-modal combinations. Specifically, the selected optimal control action sequence is denoted as ${\pi}_{t:T}$. The cost function weights are denoted as $\omega_{i}(i={1,...,N_{f}})$.
@@ III-C1 Differentiable Nonlinear Optimization
${\pi}^{*}=\arg\min_{\pi}\ \frac{1}{2}\sum_{i}^{N_{f}}\|\omega_{i}c_{i}(\pi)\|^{2}$ (5)
where $c_{i}(\pi)$ is the cost function of the control action sequence $\pi$, $\omega_{i}(i=1,...,N_{f})$ is the scaling weight and $N_{f}$ is the number of cost functions $c_{i}(\pi)$. The trajectory features are designed based on autonomous driving planning tasks, which include travel efficiency, ride comfort, lane departure, traffic rules, and most importantly safety [22].
@@ III-C2 Maximum Entropy Inverse Reinforcement Learning
$P(\xi|\omega)=\frac{\exp(-c(\xi|\omega))}{Z(\omega)}$ (6)
where $Z(\omega)$, called partition function, equals to $\sum_{\xi}\exp(-c(\xi|\omega))$. According to this function, plans with equivalent costs have equal probability.
The Max-Ent IRL method optimizes the weight $\omega$ by maximizing the probability of the human demonstration trajectories $\xi\in D$. The gradient of this optimization problem can be computed as follows [26]:
$\nabla\mathcal{L}_{IRL}=\mathbb{E}_{P(\xi|\omega)}[\hat{f}_{i}]-f_{i}$ (7)
@@ III-C3 Instance-Based Transfer Imitation Learning
The parameter $\alpha>0$ denotes the scaling factor of the $L_{IRL}$ term, utilized for adjusting the weight of the regularization component within the optimization objective. $L_{IRL}$ serves as a regularization term aimed at reducing the performance disparity across domains, which accomplishes style approximation by aligning the feature distribution expectation of the input sample with that of the user sample. In practical implementation, we utilize this regularization term along with an imitation learning term to update parameters across all models.
@@
initialize: $\omega\leftarrow$ initialize parameter via Pre-Training (Eq. 3)
@@ IV-C Ablation Study
  • NN: neural network, CF: cost function, NN&CF: joint neural network and cost function, Proportion: the proportion of expert data to batch data.

### v2

@@ III-A Problem Formulation
We assume the current time step is $t$. The agent state at the current time is represented as $\bm{s}_{t}=[x_{t},y_{t},\theta_{t},v_{t}]$, where $(x_{t},y_{t})$ is two-dimensional space coordinate, $\theta_{t}$ is heading angle, and $v_{t}$ is velocity. The trajectory $\xi_{0:t}$ is represented as a set of discrete points $\{s_{0},...,s_{t}\}$. The model predicts the future trajectory of agent $i$ for the next $T$ time steps, and it’s represented as $\hat{\xi}^{i}_{t:t+T}$. The output of the neural network is the control action sequence and denoted as ${\Pi}^{m}_{t:t+T}(m={1,...,N_{m}})$, where $N_{m}$ is the number of multi-modal combinations. Specifically, the selected optimal control action sequence is denoted as ${\pi}_{t:T}$. All learnable parameters are denoted as $\omega$, including the neural network parameters $\omega^{n}$ and the cost function weights $\omega^{c}$.
@@ III-C1 Differentiable Nonlinear Optimization
${\pi}^{*}=\arg\min_{\pi}\ \frac{1}{2}\sum_{i}^{N_{f}}\|\omega_{i}^{c}c_{i}(\pi)\|^{2}$ (5)
where $c_{i}(\pi)$ is the cost function of the control action sequence $\pi$, $\omega_{i}^{c}(i=1,...,N_{f})$ is the scaling weight and $N_{f}$ is the number of cost functions $c_{i}(\pi)$. The trajectory features are designed based on autonomous driving planning tasks, which include travel efficiency, ride comfort, lane departure, traffic rules, and most importantly safety [22].
@@ III-C2 Maximum Entropy Inverse Reinforcement Learning
$P(\xi|\omega^{c})=\frac{\exp(-c(\xi|\omega^{c}))}{Z(\omega^{c})}$ (6)
where $Z(\omega^{c})$, called partition function, equals to $\sum_{\xi}\exp(-c(\xi|\omega^{c}))$. According to this function, plans with equivalent costs have equal probability.
The Max-Ent IRL method optimizes the weight $\omega^{c}$ by maximizing the probability of the human demonstration trajectories $\xi\in D$. The gradient of this optimization problem can be computed as follows [26]:
$\nabla\mathcal{L}_{IRL}=\mathbb{E}_{P(\xi|\omega^{c})}[\hat{f}_{i}]-f_{i}$ (7)
@@ III-C3 Instance-Based Transfer Imitation Learning
The parameter $\alpha>0$ denotes the scaling factor of the $L_{IRL}$ term, utilized for adjusting the weight of the regularization component within the optimization objective. $L_{IRL}$ serves as a regularization term aimed at reducing the performance disparity across domains, which accomplishes style approximation by aligning the feature distribution expectation of the input sample with that of the user sample. In practical implementation, we utilize this regularization term along with an imitation learning term to update all learnable parameters $\omega$.
@@
initialize: $\omega^{n}\leftarrow$ initialize parameter via PreTraining (Eq. 3)
@@ IV-C Ablation Study
  • NN: neural network, CF: cost function, NN&CF: joint neural network and cost function, Proportion: the proportion of expert data in the mini-batch of input.